# Teleportation=Translation: Continuous recovery of black hole information

### v3

@@ 3.4 The Unitary Path and its Generator
The distinction between the pedagogical path (Sec. 3.2) and the physical path (Sec. 3.3) is not merely technical but conceptual. What is required is not just a continuous family of completely positive maps connecting the endpoints, but a one-parameter family of conditional expectations whose ranges remain von Neumann algebras. Only then do the modular objects exist along the entire path allowing for the rigorous definition and differentiation of the unitary flow that implements information recovery.
  • Pedagogical Path (Algebraic Limitation): The linear CP-interpolation $E_{s}=(1-s)\mathrm{id}+sE$ provides a mathematically well-defined path of maps. However, as illustrated in the $M_{2}(\mathbb{C})$ model (Appendix B), this linearity generally does not preserve the idempotency condition ($E_{s}^{2}\neq E_{s}$) for intermediate values of $s$. This property holds regardless of the algebra’s Type or the existence of a trace. As a result, the image of $E_{s}$ does not form a von Neumann subalgebra, which precludes the definition of the relative modular operator associated with the inclusion necessary to generate the unitary flow.
  • Physical Path (Algebraic Consistency): Conversely, the path $\tilde{E}_{s}$, defined by the canonical $L^{p}$-interpolation within the semifinite envelope, is idempotent by construction (Theorem 3.6). Since its image $\tilde{\mathcal{N}}(s)$ forms a von Neumann subalgebra for all $s\in[0,1]$, modular theory remains applicable throughout the interpolation. This structure ensures that the path represents a genuine flow of algebras rather than a simple interpolation of states, enabling the construction of the generating Hamiltonian even within the Type III setting.
The structural consistency of the physical path enables an explicit definition of the modular conjugation $J_{\tilde{\mathcal{N}}(s)}$ and the strongly continuous unitary path:
@@
This specific construction represents the continuous counterpart to the operational realization of the discrete information recovery process discussed in vdHV and summarized in Sec. 2.2. Applying the logic established in Step 3 of the protocol, the modular conjugation acts as a teleportation map at each $s$, relocating the instantaneous relative commutant $\tilde{\mathcal{N}}(s)^{\prime}\cap\tilde{\mathcal{M}}$ to the corresponding basic extension’s commutant $\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$, where
@@
Consequently, the flow generated by $\tilde{U}(s)$ explicitly implements this transport: it progressively decodes the information hidden in the relative commutant sector, mapping it back into the decodable sector, thereby effectively reversing the coarse-graining induced by the inclusion. Crucially, the analytic nature of the Haagerup–Kosaki interpolation guarantees the boundary conditions $\tilde{U}(0)=\mathbf{1}$ and $\tilde{U}(1)=U_{\tilde{\Gamma}}$, ensuring that the path continuously connects the identity to the target canonical shift.
The path $\tilde{U}(s)$ does not generally satisfy the group property $\tilde{U}(s+t)=\tilde{U}(s)\tilde{U}(t)$. Consequently, Stone’s Theorem does not directly apply to define a generator $\tilde{G}$ such that $\tilde{U}(s)=e^{-is\tilde{G}}$. Instead, we define $\tilde{G}$ as the infinitesimal generator (the tangent vector) at $s=0$. For this operator to be well-defined and physically meaningful, we must ensure that the path is strongly differentiable on a suitable domain and $\tilde{G}$ is essentially self-adjoint. To establish these properties, we consider a dense set of vectors $\mathcal{D}_{\text{core}}\subset\mathcal{H}$ that are analytic with respect to the modular flow $\sigma_{t}^{\tilde{\phi}_{0}}$. This choice of core allows us to control the domain of the generator and provides the necessary foundation for proving its self-adjointness in Lemma 3.10.
@@
The modular covariance of the conditional expectation $\tilde{E}_{s}$ is the pivotal property ensuring the structural consistency of this construction. Specifically, the fact that $\tilde{E}_{s}$ commutes with the modular flow $\sigma_{t}^{\tilde{\phi}_{0}}$ guarantees that analyticity is preserved under projection: analytic elements—those possessing entire modular continuations—are mapped to analytic elements within the subalgebra. Concretely, if $x\in\tilde{\mathcal{M}}$ admits an entire extension $t\mapsto\sigma_{t}^{\tilde{\phi}_{0}}(x)$, its image $\tilde{E}_{s}(x)$ likewise admits an entire extension. Consequently, the common analytic core $\mathcal{D}_{\text{core}}$ remains invariant under the interpolated expectations for all $s$. This invariance provides the stable, dense domain required to strictly define $\tilde{G}$ as a derivation, serves as the foundation for proving its self-adjointness in Lemma 3.10, and facilitates the broader analysis of its properties in Sec. 4.
The immediate physical consequence of this core preservation is that the constructed path is sufficiently smooth to define the generator $\tilde{G}$. Specifically, the invariant core $\mathcal{D}_{\text{core}}$ ensures regularity with respect to the modular flow, allowing for well-defined differentiation at $s=0$:
@@
The map $s\mapsto\tilde{U}(s)$ constructed via the canonical $L^{p}$-interpolation is real-analytic for $s\in(0,1)$. Moreover, for vectors in the common modular core $\mathcal{D}_{\text{core}}$, this map is strongly differentiable at the boundary $s=0$.
@@
    Differentiability at the Boundary: The differentiability at $s=0$ hinges on the invariance of the analytic core established in Remark 3.7. Recall that vectors $\xi\in\mathcal{D}_{\text{core}}$ are characterized by the property that their modular flow $t\mapsto\sigma_{t}^{\tilde{\phi}_{0}}(\xi)$ extends to an entire analytic function. Since the interpolation parameter $s$ corresponds to the analytic continuation of the modular group to the imaginary axis (mathematically relating $s$ to $it$), the existence of the derivative at $s=0$ is equivalent to the existence of the analytic continuation. The invariance of $\mathcal{D}_{\text{core}}$ ensures that this analytic structure is preserved along the path, guaranteeing that the limit defining the derivative exists strictly for vectors within this core.
@@
Motivated by the differentiability established in Lemma 3.8, we define the operator $\tilde{G}$ on the common modular core $\mathcal{D}_{\text{core}}$ as the tangent vector at the origin:
@@
    Symmetry: Since $\tilde{U}(s)$ is unitary for real $s$, differentiating the identity $\tilde{U}(s)^{*}\tilde{U}(s)=\mathbf{1}$ at $s=0$ yields $\tilde{U}^{\prime}(0)^{*}\tilde{U}(0)+\tilde{U}(0)^{*}\tilde{U}^{\prime}(0)=0$. Using $\tilde{U}(0)=\mathbf{1}$, this implies $\tilde{U}^{\prime}(0)^{*}=-\tilde{U}^{\prime}(0)$. Thus, $\tilde{G}=i\tilde{U}^{\prime}(0)$ is a symmetric operator ($\tilde{G}^{*}=\tilde{G}$) on the dense domain $\mathcal{D}_{\text{core}}$.
@@
    Analytic Vectors: Recall that the generator $\tilde{G}$ arises from the analytic continuation of the Radon-Nikodym cocycle (effectively the logarithm of the relative modular operator $h$). The domain $\mathcal{D}_{\text{core}}$ is explicitly defined as the set of vectors that are analytic with respect to the modular flow generated by $\tilde{\phi}_{0}$ (and by extension, $h$). Mathematically, vectors that are analytic for a one-parameter group $h^{it}$ are analytic vectors for its generator $\log h$. Since $\tilde{G}$ is linearly related to this generator at $s=0$, the set $\mathcal{D}_{\text{core}}$ constitutes a dense set of analytic vectors for $\tilde{G}$.
@@
    Conclusion: By Nelson’s analytic vector theorem, a symmetric operator possessing a dense set of analytic vectors is essentially self-adjoint. Therefore, the closure of $\tilde{G}$ defines a unique self-adjoint quantum generator.
@@ 4 DERIVATION OF THE OPERATOR IDENTITY $\tilde{G}=2P$
Having established the definition of the self-adjoint generator $\tilde{G}$ in Sec. 3.4, we now turn to the central objective of this work: establishing the identification of the teleportation generator with the modular momentum. This section presents a derivation of the identity $\tilde{G}=2P$ for general Type III algebras. We begin by analyzing the half-sided modular inclusion (HSMI) setting to motivate the specific form of the identity, particularly the factor of 2 (Sec. 4.1). Subsequently, we employ modular perturbation theory to derive the identity as an exact relation between closed operators (Sec. 4.2). We further examine the geometric stability of this result within the framework of non-commutative geometry (Sec. 4.3) and conclude by proposing a correlation function test applicable to holographic models (Sec. 4.4).
@@ 4.1 Motivation from Half-Sided Modular Inclusions
A strong theoretical foundation for the identification of the teleportation generator with modular momentum, as proposed by vdH-V vdHV , is found in highly symmetric settings, exemplified by HSMI. In this symmetric context, the connection between the canonical shift and modular operators is not merely a hypothesis but a rigorous theorem, established by Borchers and Wiesbrock Borchers92 ; Wiesbrock93 .
As discussed by vdH-V, for an HSMI $\mathcal{N}\subset\mathcal{M}$, the discrete canonical shift unitary $U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}$ is generated by twice the spacetime translation operator $P$ (which formally corresponds to the difference of modular Hamiltonians, $K_{\mathcal{M}}-K_{\mathcal{N}}$):
@@
Our canonical interpolation constructs a strongly continuous path $\tilde{U}(s)$ that interpolates from the identity to this discrete shift, i.e., $\tilde{U}(0)=\mathbf{1}$ and $\tilde{U}(1)=U_{\tilde{\Gamma}}$. Assuming that this path is generated by a scale-independent operator $\tilde{G}$ (such that $\tilde{U}(s)=e^{-is\tilde{G}}$), the endpoint condition yields:
$e^{-i\tilde{G}}=\tilde{U}(1)=U_{\tilde{\Gamma}}=e^{-2iP}.$ (21)
This exact result in the HSMI limit strongly motivates the identification $\tilde{G}=2P$ for the general case. The factor of 2 in this relationship is dictated by the underlying modular structure. Physically, the canonical shift $U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}$ is composed of two consecutive modular conjugations. Since each modular conjugation acts as a geometric reflection (analogous to a CPT inversion), combining two such reflections results in a net translation that is exactly double the elementary geometric shift. This geometric logic justifies the factor of 2 in our operator identity.
@@ 4.2 Geometric Derivation via Canonical Path
Lemma 4.1 (Linearity of the Modular Hamiltonian).
For the canonical interpolation path $\tilde{\mathcal{N}}(s)$ defined via the Connes–Takesaki cocycle derivative $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}$, the modular Hamiltonian $K(s)=K_{\tilde{\mathcal{N}}(s)}$ satisfies the following first-order perturbation relation at $s=0$:
@@
Let $\tilde{\phi}_{0}$ be a faithful normal weight on $\tilde{\mathcal{M}}$. We define the weight $\tilde{\phi}_{1}:=\tilde{\phi}_{0}\circ\tilde{E}$, thereby extending the modular dynamics of the subalgebra $\tilde{\mathcal{N}}$ to the full algebra $\tilde{\mathcal{M}}$. By identifying the modular Hamiltonians via the relation $\Delta=e^{-K}$ (specifically, $\Delta_{\tilde{\mathcal{M}}}=e^{-K_{\tilde{\mathcal{M}}}}$ and $\Delta_{\tilde{\mathcal{N}}}=e^{-K_{\tilde{\mathcal{N}}}}$), the Connes–Takesaki Radon–Nikodym cocycle is explicitly given by:
$D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{t}\;:=\;\Delta_{\tilde{\phi}_{1}}^{it}\Delta_{\tilde{\phi}_{0}}^{-it}~.$ (23)
The canonical interpolation path $\tilde{\phi}_{s}$ is then constructed via the analytic scaling of this cocycle:
$D\tilde{\phi}_{s}:D\tilde{\phi}_{0}]_{t}\;=\;[D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{\,st}~.$ (24)
Using the cocycle identity $D\tilde{\phi}_{s}:D\tilde{\phi}_{0}]_{t}=\Delta_{s}^{it}\Delta_{0}^{-it}$ (where $\Delta_{0}=\Delta_{\tilde{\mathcal{M}}}$), we can express the modular operator $\Delta_{s}$ along the path as:
$\Delta_{s}^{\,it}\;=\;\big[D\tilde{\phi}_{1}:D\tilde{\phi}_{0}\big]_{st}\,\Delta_{0}^{\,it}\,,\qquad\forall\,t\in\mathbb{R}~.$ (25)
We now compute the derivative of (25) with respect to $s$ at the origin. First, observe that the cocycle $D\tilde{\phi}_{1}:D\tilde{\phi}_{0}]_{u}$ forms a unitary group in the parameter $u=st$, generated by the relative modular Hamiltonian $K_{\tilde{\mathcal{N}}}-K_{\tilde{\mathcal{M}}}$. Applying the chain rule, the derivative of the cocycle term is:
[…]
Consequently, differentiating the full expression in (25) yields:
$\frac{d}{ds}\Delta_{s}^{\,it}\Big|_{s=0}\;=\;\left(\frac{d}{ds}\big[D\tilde{\phi}_{1}:D\tilde{\phi}_{0}\big]_{st}\Big|_{s=0}\right)\Delta_{0}^{\,it}\;=\;-it(K_{\tilde{\mathcal{N}}}-K_{\tilde{\mathcal{M}}})\Delta_{0}^{\,it}~.$ (27)
Alternatively, we can evaluate this derivative using the definition $\Delta_{s}^{it}=e^{-itK(s)}$. Applying Duhamel’s formula for the derivative of an exponential operator Kato , a cornerstone of Araki’s modular perturbation theory Araki:1976zv , we obtain:
@@
[…]
$K^{\prime}(0)\;=\;K_{\tilde{\mathcal{N}}}-K_{\tilde{\mathcal{M}}}~.$ (29)
Recalling the definition $P:=K_{\tilde{\mathcal{M}}}-K_{\tilde{\mathcal{N}}}$, we conclude that $K^{\prime}(0)=-P$. ∎
The result $K^{\prime}(0)=-P$ implies that the infinitesimal change of the modular Hamiltonian with respect to $s$ is governed by the spacetime translation operator $P$. As we vary $s$, the evolution of $K(s)$ is effectively driven by a geometric translation in the emergent spacetime. This confirms that our continuous interpolation path aligns with a physical flow: at $s=0$, the infinitesimal change of the modular Hamiltonian is precisely given by P-P (implying a translation directed opposite to the canonical outward shift). Thus, the abstract path $\tilde{E}_{s}$ is not merely algebraically idempotent; its tangent direction corresponds to a physical translation, substantiating the interpretation of the channel as a geometric transport of information.

### v4

@@ 3.4 The Unitary Path and its Generator
Leveraging the structural consistency established in Theorem 3.5, we now proceed to construct the unitary recovery operator. Since the image $\tilde{\mathcal{N}}(s)$ forms a valid von Neumann subalgebra for all $s\in[0,1]$, the associated modular objects are well-defined along the entire trajectory. This allows us to explicitly define the modular conjugation $J_{\tilde{\mathcal{N}}(s)}$ and, consequently, the strongly continuous unitary path:
@@
This construction represents the continuous, field-theoretic counterpart to the discrete information recovery protocols discussed in vdHV and summarized in Sec. 2.2. Following the logic established in Step 3 of the protocol, the operator $\tilde{U}(s)$ acts as a dynamic teleportation map. At each instant $s$, the combined action of the modular conjugations relocates the information from the instantaneous relative commutant $\tilde{\mathcal{N}}(s)^{\prime}\cap\tilde{\mathcal{M}}$ to the dual relative commutant $\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$, within the extension defined by the Jones basic construction:
@@
Physically, the flow generated by $\tilde{U}(s)$ explicitly implements this transport: it progressively decodes the information hidden in the relative commutant sector, mapping it back into the decodable sector, thereby effectively reversing the coarse-graining induced by the inclusion. Crucially, the analytic nature of the Haagerup–Kosaki interpolation guarantees the boundary conditions $\tilde{U}(0)=\mathbf{1}$ and $\tilde{U}(1)=\tilde{U}_{\Gamma}$, ensuring that the path continuously connects the identity to the target canonical shift.
The path $\tilde{U}(s)$ does not generally satisfy the group property $\tilde{U}(s+t)=\tilde{U}(s)\tilde{U}(t)$. Consequently, Stone’s Theorem does not directly apply to define a generator $\tilde{G}$ such that $\tilde{U}(s)=e^{-is\tilde{G}}$. Instead, we define $\tilde{G}$ as the infinitesimal tangent vector at the origin $s=0$. For this operator to be well-defined and physically meaningful, we must ensure that the path is strongly differentiable on a suitable domain and that $\tilde{G}$ is essentially self-adjoint. To establish these properties, we restrict our attention to a dense core of vectors $\mathcal{D}_{\text{core}}\subset\mathcal{H}$ that are analytic with respect to the modular flow $\sigma_{t}^{\tilde{\phi}_{0}}$. This choice of domain provides the necessary analytic control to prove the self-adjointness of the generator in Lemma 3.9.
@@
The structural consistency of this construction rests on the modular covariance of the conditional expectations $\tilde{E}_{s}$. Specifically, the fact that $\tilde{E}_{s}$ commutes with the modular automorphism group $\sigma_{t}^{\tilde{\phi}_{0}}$ (i.e., $\tilde{E}_{s}\circ\sigma_{t}^{\tilde{\phi}_{0}}=\sigma_{t}^{\tilde{\phi}_{0}}\circ\tilde{E}_{s}$) guarantees that analyticity is preserved under projection: elements that are analytic with respect to the modular flow are mapped to analytic elements within the subalgebra. More precisely, if $x\in\tilde{\mathcal{M}}$ admits an entire analytic continuation $t\mapsto\sigma_{t}^{\tilde{\phi}_{0}}(x)$, its image $\tilde{E}_{s}(x)$ admits the same entire extension. Consequently, the common analytic core $\mathcal{D}_{\text{core}}$ remains invariant under the interpolated expectations for all $s\in[0,1]$. This invariance provides the stable, dense domain required to rigorously define $\tilde{G}$ as a derivation and serves as the foundation for proving its self-adjointness in Lemma 3.9, and facilitates the broader analysis of its properties in Sec. 4.
The immediate physical consequence of this core preservation is that the constructed path is sufficiently smooth to define a generator. The invariant core $\mathcal{D}_{\text{core}}$ ensures regularity with respect to the modular flow, enabling well-defined differentiation at the boundary $s=0$.
@@
The map $s\mapsto\tilde{U}(s)$ constructed via the canonical $L^{p}$-interpolation is real-analytic for $s\in(0,1)$. Moreover, for vectors in the common modular core $\mathcal{D}_{\text{core}}$, the map is strongly differentiable at the boundary $s=0$.
@@
    Differentiability at the Boundary: The differentiability at $s=0$ hinges on the invariance of the analytic core established in Remark 3.6. Recall that the core $\mathcal{D}_{\text{core}}$ consists of vectors $\xi$ for which the map $t\mapsto\sigma_{t}^{\tilde{\phi}_{0}}(\xi)$ extends to an entire analytic function. As established in Definition 3.4, the parameter $s$ governs the non-commutative $L^{p}$-resolution via $s=1/p$, a structure intrinsically defined by the analytic continuation of the modular flow to the imaginary axis (Wick rotation $t\to-is$). Consequently, the derivative with respect to $s$ at the boundary ($s=0$, corresponding to $p=\infty$) is directly generated by the modular Hamiltonian. The invariance of $\mathcal{D}_{\text{core}}$ ensures that this analytic structure is preserved along the path, guaranteeing that the strong limit defining the derivative exists for all vectors within this core.
@@
Motivated by the differentiability established in Lemma 3.7, we define the operator $\tilde{G}$ on the common modular core $\mathcal{D}_{\text{core}}$ as the infinitesimal generator at the origin:
@@
    Symmetry: Since $\tilde{U}(s)$ is unitary for real $s$, differentiating the identity $\tilde{U}(s)^{*}\tilde{U}(s)=\mathbf{1}$ at $s=0$ yields $\tilde{U}^{\prime}(0)^{*}\tilde{U}(0)+\tilde{U}(0)^{*}\tilde{U}^{\prime}(0)=0$. Using the boundary condition $\tilde{U}(0)=\mathbf{1}$, this implies $\tilde{U}^{\prime}(0)^{*}=-\tilde{U}^{\prime}(0)$. Thus, $\tilde{G}=i\tilde{U}^{\prime}(0)$ is a symmetric operator ($\tilde{G}^{*}=\tilde{G}$) on the dense domain $\mathcal{D}_{\text{core}}$.
@@
    Analytic Vectors: The generator $\tilde{G}$ is constructed from the analytic continuation of the Radon-Nikodym cocycle, which is generated by the logarithm of the relative modular operator $h$. The domain $\mathcal{D}_{\text{core}}$ is explicitly defined as the set of vectors that are analytic with respect to the modular flow $\sigma_{t}^{\tilde{\phi}_{0}}$ generated by $\tilde{\phi}_{0}$ (and consequently the unitary group $h^{it}$). According to the general theory of one-parameter unitary groups, the set of analytic vectors for the group $h^{it}$ constitutes a dense set of analytic vectors for its generator $\log h$. Since the action of $\tilde{G}$ at $s=0$ is linearly related to this generator, the set $\mathcal{D}_{\text{core}}$ forms a dense set of analytic vectors for $\tilde{G}$ as well.
@@
    Conclusion: By Nelson’s analytic vector theorem, a symmetric operator that admits a dense set of analytic vectors is essentially self-adjoint. Consequently, the closure $\overline{\tilde{G}}$ is self-adjoint and unique. This guarantees that $\tilde{G}$ unambiguously defines a valid quantum observable, providing the rigorous basis for the infinitesimal generation of the recovery flow.
@@
Remark 3.11 (Canonical Nature and Covariance).
The construction of the generator $\tilde{G}$ is free from the ambiguities typically associated with cutoff-based regularization schemes. The canonical $L^{p}$-interpolation path is uniquely determined solely by the initial reference weight $\tilde{\phi}_{0}$ and the algebraic inclusion $\tilde{\mathcal{N}}\subset\tilde{\mathcal{M}}$, independent of arbitrary choices such as basis vectors or auxiliary projections. Consequently, $\tilde{G}$ is not merely a mathematical artifact but a canonical object of the theory: it acts as an intrinsic physical observable associated with the inclusion pair, transforming covariantly under automorphisms that preserve this structure. This establishes $\tilde{G}$ as the rigorously defined, self-adjoint generator of the continuous teleportation protocol.
@@ 4 DERIVATION OF THE OPERATOR IDENTITY $\tilde{G}=2P$
With the self-adjoint generator $\tilde{G}$ rigorously defined in Sec. 3.4, we now address the central objective of this work: identifying the teleportation generator with the modular momentum. This section presents a derivation of the identity $\tilde{G}=2P$ for general Type III algebras. We begin by motivating the specific form of the identity, particularly the factor of 2, through the limit of HSMI (Sec. 4.1). We then employ modular perturbation theory to prove the identity as an exact relation between closed operators (Sec. 4.2). Finally, we examine the geometric stability of this result within the framework of non-commutative geometry (Sec. 4.3) and propose a holographic correlation function test to verify its physical validity (Sec. 4.4).
@@ 4.1 Motivation from Half-Sided Modular Inclusions
A strong theoretical foundation for identifying the teleportation generator with modular momentum, as proposed by vdH-V vdHV , is found in highly symmetric settings, exemplified by HSMI. In this context, the connection between the canonical shift and modular operators is not merely a hypothesis but a rigorous theorem established by Borchers and Wiesbrock Borchers92 ; Wiesbrock93 .
As discussed in vdHV , for an HSMI $\mathcal{N}\subset\mathcal{M}$, the discrete canonical shift unitary $U_{\Gamma}=J_{\mathcal{M}}J_{\mathcal{N}}$ is generated by exactly twice the spacetime translation operator $P$ corresponding to the difference of modular Hamiltonians, $K_{\mathcal{M}}-K_{\mathcal{N}}$.
@@
Our canonical interpolation provides a dynamic generalization of this structure. Recall that we constructed a differentiable unitary path $\tilde{U}(s)$ satisfying the boundary conditions $\tilde{U}(0)=\mathbf{1}$ and $\tilde{U}(1)=U_{\tilde{\Gamma}}$. While this path does not generally form a one-parameter group in generic Type III settings, in the strict HSMI limit, the path is known to take the exact exponential form $\tilde{U}(s)=e^{-2isP}$. Consequently, applying the definition of the generator $\tilde{G}:=i\tilde{U}^{\prime}(0)$ given in Eq. (20) directly yields:
$\tilde{G}=i\frac{d}{ds}\left(e^{-2isP}\right)\bigg|_{s=0}=2P.$ (22)
This derivation provides a geometric motivation for the factor of 2. Physically, $\tilde{U}(s)=J_{\tilde{\mathcal{M}}}J_{\tilde{\mathcal{N}}(s)}$ represents the composition of a fixed modular reflection ($J_{\tilde{\mathcal{M}}}$) and one shifted by the parameter $s$ ($J_{\tilde{\mathcal{N}}(s)}$). Analogous to Euclidean geometry, where composing two reflections separated by a distance $d$ generates a translation of $2d$ (which corresponds to the full teleportation limit at $s=1$), this composite action induces a shift of twice the modular parameter. This geometric kinematics directly implies the generator relation $\tilde{G}=2P$.
@@ 4.2 Geometric Derivation via Canonical Path
Lemma 4.1 (First-Order Variation of the Modular Hamiltonian).
For the canonical interpolation path $\tilde{\mathcal{N}}(s)$ defined in Definition 3.4, the modular Hamiltonian $K(s)=K_{\tilde{\mathcal{N}}(s)}$ satisfies the following perturbation relation at the origin $s=0$:
@@
We utilize the structural definitions established in Definition 3.4. Recall that the interpolation path is generated by the analytic continuation of the Connes-Takesaki cocycle associated with the Radon-Nikodym derivative $h=d\tilde{\phi}_{1}/d\tilde{\phi}_{0}$. The interpolated cocycle satisfies the scaling relation:
$D\tilde{\phi}_{s}:D\tilde{\phi}_{0}]_{t}=h^{ist}~.$ (24)
The Connes cocycle intertwines the modular operators via the identity
$\Delta_{s}^{it}=[D\tilde{\phi}_{s}:D\tilde{\phi}_{0}]_{t}\Delta_{0}^{it}~,$ (25)
where $\Delta_{s}=e^{-K(s)}$ and $\Delta_{0}=e^{-K_{\tilde{\mathcal{M}}}}$. Combining this with Eq. (24), we express the modular operator along the path directly in terms of the generator $h$:
$\Delta_{s}^{it}=h^{ist}\Delta_{0}^{it},\qquad\forall\,t\in\mathbb{R}~.$ (26)
We now compute the derivative with respect to $s$ at $s=0$. Differentiating the right-hand side of (26) yields:
$\frac{d}{ds}\left(h^{ist}\Delta_{0}^{it}\right)\Big|_{s=0}=it(\ln h)\Delta_{0}^{it}~.$ (27)
To identify the operator $\ln h$, we examine the generator of the cocycle at the endpoint $s=1$. Differentiating the relation $h^{it}=\Delta_{1}^{it}\Delta_{0}^{-it}$ with respect to $t$ at $t=0$ yields $i\ln h$ on the left-hand side. On the right-hand side, the derivative gives $(-iK_{\tilde{\mathcal{N}}})-(-iK_{\tilde{\mathcal{M}}})=i(K_{\tilde{\mathcal{M}}}-K_{\tilde{\mathcal{N}}})=iP$. Comparing these generators establishes the identity $\ln h=P$. Consequently, the derivative in (27) simplifies to $itP\Delta_{0}^{it}$.
Alternatively, we evaluate the derivative of the left-hand side of (26), explicitly writing $\Delta_{s}^{it}=e^{-itK(s)}$. Applying Duhamel’s formula for the derivative of an exponential operator Kato , a cornerstone of Araki’s modular perturbation theory Araki:1976zv , we obtain:
@@
To extract the generator $K^{\prime}(0)$, we divide both sides by $t$ and take the limit $t\to 0$. In this limit, the integral term converges to $iK^{\prime}(0)$ due to the strong continuity of the modular flow. Specifically, continuity ensures that $e^{-i\tau K_{0}}\to\mathbf{1}$ strongly as $\tau\to 0$, rendering non-commutative effects negligible at the leading order. On the other hand, for the cocycle term (the right-hand side), we have $\lim_{t\to 0}\frac{1}{t}(itP\Delta_{0}^{it})=iP$ (since $\Delta_{0}^{it}\to\mathbf{1}$). Comparing these limits establishes the identity:
$iK^{\prime}(0)=iP\quad\implies\quad K^{\prime}(0)=-P~.$ (29)
This confirms that the infinitesimal variation of the modular Hamiltonian corresponds exactly to the negative of the modular momentum. ∎
The result $K^{\prime}(0)=-P$ implies that the infinitesimal change of the modular Hamiltonian with respect to $s$ is governed by the spacetime translation operator $P$. As the interpolation parameter $s$ varies, the evolution of $K(s)$ is effectively driven by a geometric translation within the emergent spacetime. This confirms that our canonical interpolation path aligns with a distinct physical flow: at the boundary $s=0$, the infinitesimal change of the modular Hamiltonian is precisely given by P-P, indicating a trajectory directed towards the subalgebra (opposite to the canonical outward growth). Consequently, the path of conditional expectations $\tilde{E}_{s}$ is not merely a mathematical abstraction of subalgebra inclusion; its tangent direction manifests as a concrete physical translation. This substantiates the interpretation of the channel as a geometric transport of information, providing the dynamical basis for the recovery protocol.